OpenStudy (anonymous):

Find the average value of f(x) = sin(x) over the interval [0, π]. a: 2 b: 2π c: 2/ pi d: 0

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@freckles

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Abhisar

OpenStudy (loser66):

I don't know :(

OpenStudy (misty1212):

hi!!

OpenStudy (anonymous):

its ok, thank u anyways!! :)

OpenStudy (anonymous):

@misty1212 hi :)

OpenStudy (misty1212):

\[\frac{1}{\pi}\int _0^{\pi}\sin(x)dx\] is the average value

OpenStudy (misty1212):

i wouldn't bother computing the integral

OpenStudy (anonymous):

Calculating integrals :(

OpenStudy (anonymous):

kk

OpenStudy (misty1212):

only one answer, as usual C, has a \(\pi\) in the denominator

OpenStudy (anonymous):

ur so awesome!!!!!!!!!!!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]